OpenStudy (anonymous):

Simplify 1 · x - x/1 0 1 x

OpenStudy (anonymous):

Do you mean,\[\frac{x-x}{1}\]Or,\[x-\frac{x}{1}\]?

OpenStudy (anonymous):

the 2nd one

OpenStudy (anonymous):

\[x-\frac{x}{1}=\frac{x}{1}-\frac{x}{1}=x-x=0\]

OpenStudy (anonymous):

thank you.

OpenStudy (anonymous):

No problem :)